OpenStudy (anonymous):

The length of a rectangle is 5 units and its width is 4 units. What is the approximate length of the diagonal of the rectangle? 5 units 6.4 units 8.5 units 9 units

OpenStudy (anonymous):

9 I believe

OpenStudy (anonymous):

I NEED HELP ON THIS ALSO

OpenStudy (er.mohd.amir):

diagonal=sqrt (length^2+widthh^2)

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

Yeah. That would be 9.

OpenStudy (ddcamp):

@asyl98 Do you know the Pythagorean Theorem?

OpenStudy (anonymous):

thank u

OpenStudy (er.mohd.amir):

D*D=5*5+4*4=25+16=41 find square root of 41 is ans

OpenStudy (ddcamp):

The answer is definitely not 9...